# Emergent Hydrodynamic Mode on SU(2) Plaquette Chains and Quantum Simulation

### v1

@@ Appendix A Time to reach boundary
To study the boundary effect, we estimate the time it takes for a physical signal to reach the boundary. In particular, we calculate the symmetric correlator $G_{s}(t,j,i)$ at two different positions $j=0$ and $i=8$ on an open $N=17$ plaquette chain with $ag^{2}=1.2$ and $j_{\rm max}=\frac{1}{2}$. Results at two different temperatures $T=5$ and $T=20$ are shown in Fig. 6. We observe that the perturbation introduced at the center of the lattice $i=8$ at $t=0$ takes about $t=12.5$ to reach the boundary, at which it bounces back on the open lattice. It is estimated that the perturbation bounced back from the boundary reaches the center of the lattice at roughly $t=25$.
@@ Appendix B Symmetric correlator in microcanonical ensemble
Table 5: Parameter values of Eq. (43) fitted from the $G_{s}^{(mc)}(omega,k)$ results in Fig. 7. Numbers are rounded to three significant figures.

### v2

@@ Appendix A Time to reach boundary
To study the boundary effect, we estimate the time it takes for a physical signal to reach the boundary. In particular, we calculate the symmetric correlator $G_{s}(t,j,i)$ at two different positions $j=0$ and $i=8$ on an open $N=17$ plaquette chain with $ag^{2}=1.2$ and $j_{\rm max}=\frac{1}{2}$. Results at two different temperatures $T=5$ and $T=20$ are shown in Fig. 6. We observe that the perturbation introduced at the center of the lattice $i=8$ at $t=0$ takes about $t=12.5$ to reach the boundary, at which point it bounces back on the open lattice. It is estimated that the perturbation bounced back from the boundary reaches the center of the lattice at roughly $t=25$.
@@ Appendix B Symmetric correlator in microcanonical ensemble
Table 5: Parameter values of Eq. (43) fitted from the $G_{s}^{(mc)}(\omega,k)$ results in Fig. 7. Numbers are rounded to three significant figures.